OpenStudy (nuccioreggie):

Which of the following is a rational number? square root 97square root 98, square root 99, square root 100 square root 97 square root 98 square root 99 square root 100

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

97 is a prime number 98= 7*7*2, which is not a perfect square: 99=3*3*11 which is not a perfect square 100=10*10 which is a perfect square